# Fundamental mechanisms of high-electric-field vacuum arc initiation

### v1

@@ III.3 Breakdown Statistics
We now consider the statistical characteristics of breakdowns which emerge when considering an ensemble of breakdowns. One of the most directly observable and important statistical properties of a pulsed high-field device is that it operates with a breakdown rate. The breakdown rate is defined as the number of breakdowns divided by the number of pulses during a certain measurement window. Stable operation of a large-scale facility like the CLIC linear collider creates the requirement that breakdown rate of individual components, in particular the accelerating structures, be very low . The dependencies of breakdown rate on parameters such as field level, pulse length and temperature are very strong and will be discussed in depth in Sec. III.5.
@@
When a high field device operates for an extended period, a distribution of intervals between breakdowns emerges that can be visualized by a histogram of the number of pulses between breakdowns. A typical example of such histograms in dc and radio frequency are shown in Fig.13. The histogram is not that of a simple Poisson distribution, which would be the case if breakdowns occurred randomly, but rather has two parts, each Poisson-like. The longer interval part is caused by the underlying, long-term breakdown rate. The shorter interval part, occupying typically the first few thousand pulses is a caused by breakdowns triggered by previous breakdown. Breakdowns in the two parts of the distribution are referred to as primary and follow-up breakdowns. More data and analysis of this phenomenon can be found in [145, 185, 150, 115].
@@
Breakdowns also exhibit a spatial distribution, and we shall see that it is related to the temporal distribution. The localization of breakdowns in the pulsed dc system is made using triangulation of emitted breakdown light as we have seen in SubSec. II.3. The longitudinal position of breakdown positions in an traveling wave radio frequency structure can also be determined. The technique measures the breakdown position from the relative timing of the rising edge of the reflected wave and falling edge of the transmitted wave with respect to the beginning of the incident pulse along with the group velocity in the structure. This measurement is feasible due to the steeply rising/falling edges of the reflected/transmitted wave.
@@
Complementary position information is contained in the relative phase between the incident and the reflected wave from the breakdown. A traveling wave accelerating structure has a characteristic phase advance. One can see this phase advance in plots of the reflected waves which are grouped into distinct phases that correspond to two times the cell-to-cell phase advance in the structure. This indicates that the breakdowns and associated short circuit are localized to the same location in each cell. This is consistent with post-test imaging, as we will see below, that show breakdown craters are concentrated on the coupling iris in this type of structure. Combining timing and reflected phase gives an enhanced accuracy in determining the longitudinal position in a traveling wave accelerating structure as shown in 14. Determination of breakdown position in a traveling wave accelerating structure technique is described in more depth in [145].
@@
Temporal and spatial distributions can now be combined into a single plot, in order to motivate into the physical explanation for the two part Poissonian distribution seen in 13. Such a combined plot produced from data taken in high power testing of a traveling wave accelerating structure is shown in 15. The plot shows that that successive breakdowns separated by relatively fewer pulses tend to occur closer together, and conversely those separated by a large number of pulses tend to be further apart. The most plausible explanation for the short timescale effect is that breakdowns produce surface features, for example crater-like bumps and splashes of metal as we will see in the post-test microscopy discussion below, that result in an enhanced propensity to breakdown. The enhanced tendency to break down decreases with time, either because features are eliminated during the subsequent breakdowns or because the features are smoothed without breakdown over a certain number of pulses. Once the aftereffects of a breakdown are neutralized, the high-field surfaces return to their underlying breakdown rate and are due to the mechanisms described in section IV.
@@
From these observations, the concept of primary and follow up breakdowns has been introduced. The number of follow up breakdowns can be influenced by how fields are adjusted following a breakdown. In the many high field systems, such as radio frequency accelerating structures, power, and consequently field, levels are reduced immediately following breakdown and are ramped up in pre-defined function following breakdown. A comparative study of different post breakdown ramps was carried out in [151, 150] with the objective to find a ramp which minimized the number of follow up breakdowns.
@@
Another important statistical behavior that is observed during pulsed mode operation, is the time-to-breakdown within the pulses. An example of such a distribution is shown in Fig. 16. The striking feature of this distribution is that it is relatively flat. It is not, for example, rising towards then end of the pulse which might be the case if the breakdown initiation process was the result of a intra-pulse heating process. Further time-to-breakdown data can be found in [108, 33]. The time-to-breakdown within pulses is discussed further in Sec. IV.5.
@@ III.4 The conditioning process
Understanding what is actually changing in a device when it progresses from an unconditioned to a conditioned state has significant potential benefits both theoretical, as important input for the understanding the different mechanisms that initiate breakdown, and practical, for guiding the development of fabrication methods that reduce conditioning time.
@@
The overall shape of the conditioning curve, that of an initially steep rise followed by an asymptotic approach to the maximum stable field holding field is a typical one. However the details of the curve can be affected by the series of specific decisions including about how quickly to raise the field, how to recover stable operation following a breakdown and when to raise pulse length. We will refer this collective set of decisions as the conditioning strategy. All of the RF structures, including that shown in Fig. 17, and dc electrodes referred to in this report have been conditioned using conditioning strategies that have been implemented as computer controlled algorithms. This algorithmic control is essential for comparing results from different structures and electrodes in a quantitative way and also for producing reproducible device performances. In many domains of high-field systems conditioning is controlled by a human operator, so cannot be precisely defined, which makes comparison of results difficult.
@@
The initial stages of conditioning, when fields are applied but before the first breakdown is observed, is characterized by an increased vacuum level. This is caused by electron stimulated desorption (ESD) outgassing of the structure surface by the applied fields, typically caused by field-emitted electron bombardment of the inner surface of the cavity, emitting also UV and X-Rays photons resulting indirectly in additional photon stimulated desorption. With time, the vacuum rise seen when fields are applied decreases. The static, no-field, vacuum level seen after a period of applied field operation is lower than before the applied field period. This is because the high-field operation has an effect similar to a vacuum bake-out for the surfaces exposed to the direct and indirect effects of high fields.
@@
The vacuum behavior raises the interesting question of the role of residual gas in breakdown initiation itself. No significant trigger effect from these gases, such as a larger release prior to a breakdown, has been observed [34]. In fact, the amount of ambient gases surrounding the electrodes seems to have only a limited effect on the breakdown triggering. Experiments performed by injecting different gases (Ar, H2, CO and air) demonstrated that pressure in excess of are needed in order to lower the breakdown field from its typical saturated value [146]. This pressure value is orders of magnitude larger than the background vacuum level of the test system as mentioned earlier, and in fact also of the vacuum level of a typical radio frequency accelerating structure. It can thus be excluded than in normal operation the vacuum level has a significant influence on the breakdown triggering mechanism, at least in this type of test system. It is anyhow as yet unclear whether the observed breakdown field decrease is due to enhanced ionisation of the ambient gas by field emitted electrons, which seems unlikely due to the mean free path being orders of magnitude larger that the gap distance, or to enhanced surface adsorption of gases, and their subsequent release upon applying electric field playing a role as an early trigger for breakdowns.
@@
We will now refer to Fig. 17 to describe the subsequent stages of conditioning. After the initial vacuum-limited operation, the vacuum level will no longer rise to critical values with fields applied, so the field can be raised to a level where the breakdowns start to occur. In Fig. 17 the first breakdowns of this type occur at an accelerating gradient in the range of 10 to 30 MV/m, which corresponds to a peak surface electric field of about 25 to 75 MV/m for this type of structure. In this stage of conditioning the breakdowns produce large vacuum bursts. It is believed that these initial breakdowns are initiated primarily by extrinsic factors, such as residual dust, which contain a significant amount of non-metal material that is vaporized by the breakdown. Like the initial steady state vacuum pressure decreases during operation, the magnitude of the pressure bursts caused by the breakdowns decreases with continued operation. This transition is probably because the earliest breakdowns are caused by extrinsic factors, while the later, and vast majority in a high-field system, breakdowns are caused by the intrinsic mechanisms. Because the intrinsic mechanisms only involve copper, and copper will immediately stick to the first surface it encounters, and will not propagate to the vacuum gauge.
@@
A comparison of conditioning data from different structures of the same type leads to the observation that progress in conditioning proceeds with the number of pulses, not the number of breakdowns. This observation is illustrated in the plot in Fig. 18 which is taken from [31]. The similarity of the conditioning curves of different structures of the same type when plotted against the number of pulses, and dissimilarity when plotted against the number of breakdowns points to conditioning being driven by the number of pulses. This argues against the supposition that conditioning process is the removal by breakdown of extrinsic elements on the device surface. It argues rather for that an applied surface field drives a change to a device that both improves the surface, although simultaneously and giving rise to a continuing creation of breakdown nucleation points.
The breakdown dominated phase of conditioning is typically initially carried out at a reduced pulse length. This is done partly to minimize damage to the device by reducing the pulse energy. It is also done because it appears to decrease the conditioning time. This is because the conditioning effect of a pulse on the surface is greater at a higher field level, we will return to this point below, while at the same time because the breakdown rate is lower at shorter pulse lengths as discussed in Subsec. III.5. Thus at a lower pulse length the field level can be ramped up at a higher rate. Pulse lengths are usually increased step-wise after the maximum field has been approached. When a pulse length increase is implemented, it is necessary to reduce the field in the device to maintain the target breakdown rate. We will discuss the dependency of breakdown rate on gradient in subsection III.5. The field is then ramped up at the new pulse length, as before. This increasing pulse stage of conditioning can be seen in the drop-then-ramp structure in Fig. 17 from about 100 to 300 million pulses.
The ultimate gradient of a particular type of structure, is determined by both its design and by the techniques that were used in its fabrication. The final stage of conditioning must be done very carefully because pushing the field too close to the ultimate value can result in irreversible damage. This damage is observed as it forces the gradient to reduced significantly, as much as 20%, in order to reestablish operation at the typical target breakdown rates of to . In some cases certain materials will operate for some time at a high gradient, but eventually crash as discussed in [137]. Determining the maximum practical operating field with appropriate safety margin can be determined typically only after iterations on a device type.
@@
In addition to exposure to air, a smaller but similar effect is observed when the structure is idle for an extended period. Although the vacuum may be very good in a high-field system, there is still a finite presence of gas. While the high field surfaces become extremely clean due to the action of the applied fields, the rest of the vacuum system surface has the adsorbed gas associated with the base vacuum pressure of the system. During an idle period, gas is transferred from the non high-field regions to the high-field ones. Extrinsic-associated vacuum and breakdown activity occur once fields are applied again. This is the same effect which is proposed to explain the dependency of breakdown rate on repetition rate and pause time that will be discussed Subsec. III.5.
@@
A few important insights can be made from these observations from conditioning. One is that there is not a unique type of breakdown trigger, but rather there are those associated with extrinsic surface contaminants and those associated with intrinsic features, those coming from the fundamental properties of the device material surface or bulk. The conditioning of intrinsic features takes many pulse, hundreds of millions of pulses. Intrinsic feature conditioning is preserved even when a device is exposed to air. The success of the dislocation model in explaining the dependency of breakdown rate on field and temperature as described in Sec. IV, leads to the supposition that an evolving dislocation structure is the mechanism that underlies conditioning. However differences in dislocation structure between high-field and unexposed regions has so far not been observed directly.
@@
Different materials, and materials in different metallurgical states condition in different ways and to different levels. For example, hard copper conditions faster than soft copper as shown in Fig. 19.This supports that long-term conditioning is driven by intrinsic properties of the system. We will return to dependence on material in Subsec. III.6.
@@
A crucial question for optimizing a conditioning strategy is if the conditioning speed is dependent on the level of applied field. As we have seen conditioning is correlated to the number of pulses as shown in Fig. 18 but the overall rate may be higher with applied field. Evidence that conditioning proceeds more quickly with higher applied field is reported in [81], with key evidence reproduced here in Fig. 18. The anodes and cathode in this in the test had the same diameter, which resulted in a field enhancement at the outer edge of about 7.5%. For the hard electrode, on the left in Fig. 18, breakdowns populate the outer edge because the hard electrode is in a more conditioned state, as can be seen in the faster conditioning shown in Fig. 19. On the other hand, breakdowns populate the soft electrode quite uniformly, and are not concentrated on the outer edge despite the field enhancement. The explanation for this that the higher field on the outer edges results in a higher conditioning speed compared to the center of the electrode, which offsets the strong dependency of breakdown rate on applied field.
@@
Taking this observation on the interplay between conditioning speed and breakdown rate through applied field further, and in order to better determine the mechanisms involved in conditioning a Monte Carlo simulation tool has been developed [114], also inspired by previous work [95] . The dual role of applied field in both driving conditioning and the breakdown rate is a key element of understanding the breakdown process.

### v2

@@ III.3 Breakdown Statistics
We now consider the statistical characteristics of breakdowns which emerge when considering an ensemble of breakdowns. One of the most directly observable and important statistical properties of a pulsed high-field device is that it operates with a breakdown rate. Stable operation of a large-scale facility like the CLIC linear collider creates the requirement that breakdown rate of individual components, in particular the accelerating structures, be very low . The origin of breakdown rate will emerge from the dislocation model described in Sec. IV Plastic Material Response.
@@
When a high field device operates for an extended period, a distribution of intervals between successive breakdowns emerges that can be visualized by a histogram of the number of pulses between breakdowns. Typical examples of such histograms, one dc and one radio frequency, are shown in Fig. 14. The histograms are not simple Poisson distributions, which would be the case if breakdowns occurred randomly, but rather have two parts, each Poisson-like. Breakdowns occurring after a long interval are caused by the underlying, long-term breakdown rate mechanism and are referred to as primary breakdowns. Those occurring after a short interval, typically a few thousand pulses, are caused by the after-affects of a previous breakdown and are referred to as follow-up breakdowns. The source of a follow-up breakdown can be the droplets or sharper features seen in breakdown craters as shown described in Sec. III.8. After the most unstable features either initiate a follow-up breakdown or are eroded by field exposure, the system returns to the long-term breakdown rate. The long-term breakdown rate is given by dislocation dynamics, as will be described in Sec. IV Plastic Material Response. More data and analysis of this phenomenon can be found in [154, 195, 159, 124].
@@
There is also a distribution of the distance between successive breakdown locations, and we shall see that it is related to the distribution of the number of pulses. The localization of breakdowns in the pulsed dc system is made using triangulation of emitted breakdown light as we have seen subsection II.3. The longitudinal position of breakdown positions in a traveling wave radio frequency structure can also be determined. The technique measures the breakdown position from the relative timing of the rising edge of the reflected wave and falling edge of the transmitted wave with respect to the beginning of the incident pulse along with the group velocity in the structure. This measurement is feasible due to the steeply rising/falling edges of the reflected/transmitted wave.
@@
Complementary position information is contained in the relative phase between the incident and the reflected wave from the breakdown. A traveling wave accelerating structure has a characteristic phase advance. One can see this phase advance in plots of the reflected waves, which are grouped into distinct phases that correspond to two times the cell-to-cell phase advance in the structure. This indicates that the breakdowns and associated short circuits are localized to the same location in each cell. This is consistent with post-test imaging, as we will see below, that show that breakdown craters are concentrated on the coupling iris in this type of structure. Combining timing and reflected phase gives an enhanced accuracy in determining the longitudinal position in a traveling wave accelerating structure, as shown in Fig. 15. Determination of breakdown position in a traveling wave accelerating structure technique is described in more depth in [154].
@@
Temporal and spatial distributions can now be combined into a single plot, in order to motivate into the physical explanation for the two part Poissonian distribution seen in Fig. 14. Such a combined plot produced from data taken in high power testing of a traveling wave accelerating structure is shown in Fig. 16. The plot shows that successive breakdowns separated by relatively fewer pulses tend to occur closer together, and conversely those separated by a large number of pulses tend to be further apart. The most plausible explanation for the short timescale effect is that breakdowns produce surface features, for example crater-like bumps and splashes of metal as we will see in the post-test microscopy discussion below, that result in an enhanced propensity to breakdown. The enhanced tendency to break down decreases with time, either because features are eliminated during the subsequent breakdowns or because the features are smoothed without breakdown over a certain number of pulses. Once the aftereffects of a breakdown are neutralized, the high-field surfaces return to their underlying breakdown rate which is due to the mechanisms described in section IV.
@@
The likelihood of having a follow-up breakdown is influenced by how fields are adjusted following a breakdown. In many systems the field level is reduced on the pulse that follows immediately after a breakdown. The field is then ramped back up to the niminal operating value in a pre-defined sequence on subsequent pulses. Such a breakdown-recovery ramp typically takes a hundred to a thousand pulses in the pulsed systems described in this review. A comparative study of different post breakdown ramps, was carried out in [160, 159] with the objective to find a ramp which minimized the number of follow-up breakdowns.
@@
Another important statistical behavior that is observed during pulsed mode operation, is the time-to-breakdown within the pulses. An example of such a distribution is shown in Fig. 17. The striking feature of this distribution is that it is relatively flat, for all the different pulse lengths and is not, for example, rising towards then end of the pulse. This would be the case if breakdown initiation was the result of a heating process occurring within individual pulses. The observed overall dependence is clearly not produced by a higher density of breakdowns at the end of the pulse. This means that pulse length dependence is produced by preceding pulses, which which we can refer to as a pulse-to-pulse memory effect. Further time-to-breakdown data can be found in [116, 36]. The time-to-breakdown within pulses is discussed further in Sec. IV.5.
@@ III.4 The conditioning process
Understanding what is actually changing in a device as it progresses from an unconditioned to a conditioned state has benefits that are both theoretical, important input for the understanding the different mechanisms that initiate breakdown, and practical, for guiding the development of fabrication methods that reduce conditioning time.
@@
The overall shape of the conditioning curve, that of an initially steep rise followed by an asymptotic approach to the maximum stable field holding field is a typical one. However the details of the curve can be affected by the series of specific decisions including about how quickly to raise the field, how to recover stable operation following a breakdown and when to raise pulse length. We will refer this collective set of decisions as the conditioning strategy. All of the rf structures, including that shown in Fig. 18, and dc electrodes referred to in this report have been conditioned using conditioning strategies that have been implemented as computer controlled algorithms. This algorithmic control is essential for comparing results from different structures and electrodes in a quantitative way and also for producing reproducible device performances. In many domains of high-field systems, conditioning is controlled by a human operator, so cannot be precisely defined, which makes comparison of results difficult.
@@
The initial stages of conditioning, when fields are applied but before the first breakdown is observed, is characterized by an increased vacuum level. This is caused by electron stimulated desorption (ESD) outgassing of the structure surface by the applied fields, typically caused by field-emitted electron bombardment of the inner surface of the cavity, emitting also UV and X-Rays photons resulting indirectly in additional photon stimulated desorption. The applied field induced vacuum rise decreases as conditioning progresses as well as the field-off base vacuum level. This is because the high-field operation has an effect similar to a vacuum bake-out for the surfaces exposed to the direct and indirect effects of high fields.
@@
The vacuum behavior raises the interesting question of the role of residual gas in breakdown initiation itself. No significant trigger effect from these gases, such as a larger release prior to a breakdown, has been observed [37]. In fact, the amount of ambient gases surrounding the electrodes seems to have only a limited effect on the breakdown triggering. Experiments performed by injecting different gases (Ar, H2, CO and air) demonstrated that pressure in excess of are needed in order to lower the breakdown field from its typical saturated value [155]. This pressure value is orders of magnitude larger than the background vacuum level of the test system as mentioned earlier and also of the vacuum level of a typical radio frequency accelerating structure. Consequently ambient gasses do not appear to have a direct role in breakdown triggering. It is as yet unclear whether the observed breakdown field decrease is due to enhanced ionisation of the ambient gas by field emitted electrons, which seems unlikely due to the mean free path being orders of magnitude larger than the gap distance, or to enhanced surface adsorption of gases, and their subsequent release upon applying electric field playing a role as an early trigger for breakdowns.
@@
We will now refer to Fig. 18 to describe the subsequent stages of conditioning. After the initial vacuum-limited operation, the vacuum level will no longer rise to critical values with fields applied, so the field can be raised to a level where the breakdowns start to occur. In Fig. 18 the first breakdowns of this type occur at an accelerating gradient in the range of 10 to 30 MV/m, which corresponds to a peak surface electric field of about 25 to 75 MV/m for this type of structure. In this stage of conditioning the breakdowns produce large vacuum bursts, i.e. abrupt increases in the pressure. It is believed that these initial breakdowns are initiated primarily by extrinsic factors, such as residual dust, which contain a significant amount of non-metal material that is vaporized by the breakdown. Like the initial steady state vacuum pressure decreases during operation, the magnitude of the pressure bursts caused by the breakdowns decreases with continued operation. This transition is probably because the earliest breakdowns are caused by extrinsic factors, while the later, and vast majority in a high-field system, breakdowns are caused by the intrinsic mechanisms. Because the intrinsic mechanisms only involve copper, and copper will immediately stick to the first surface it encounters and will not propagate to the vacuum gauge.
@@
A comparison of conditioning data from different structures of the same type leads to the observation that progress in conditioning proceeds with the number of pulses, not the number of breakdowns. This observation is illustrated in the plot in Fig. 19 which is taken from [34]. The similarity of the conditioning curves of different structures of the same type when plotted against the number of pulses - and their dissimilarity when plotted against the number of breakdowns - suggests that conditioning is primarily driven by the number of pulses. This argues against the supposition that conditioning progresses through the removal by breakdown of extrinsic elements on the device surface. It argues rather that the applied surface field improves the surface, while simultaneously creating breakdown nucleation points.
The breakdown dominated phase of conditioning is initially carried out typically at a reduced pulse length. This is done partly to minimize damage to the device by reducing the pulse energy. It is also done because it appears to decrease the conditioning time. This is because the conditioning effect of a pulse on the surface is greater at a higher field level, we will return to this point below, and the breakdown rate is lower at shorter pulse lengths as discussed in Sec. III.5. Thus at a lower pulse length the field level can be ramped up at a higher rate. Pulse lengths are usually increased step-wise after the maximum field has been approached. When a pulse length increase is implemented, it is necessary to reduce the field in the device to maintain the target breakdown rate. We will discuss the dependency of breakdown rate on gradient in subsection III.5. The field is then ramped up at the new pulse length, as before. This increasing pulse stage of conditioning can be seen in the drop-then-ramp structure in Fig. 18 from about 50 to 150 million pulses.
The ultimate gradient of a particular type of structure, is determined by both its design and by the techniques that were used in its fabrication. The final stage of conditioning must be done very carefully because pushing the field too close to the ultimate value can result in irreversible damage. This damage is observed as it forces the gradient to reduced significantly, as much as 20%, in order to reestablish operation at the typical target breakdown rates of to . In some cases certain materials will operate for some time at a high gradient, but eventually crash as discussed in [146]. The maximum practical operating field, including an appropriate safety margin, can typically be established only after iterative testing on a given device type.
@@
In addition to exposure to air, a smaller but similar effect is observed when the structure is idle for an extended period. Although the vacuum may be very good in a high-field system, there is still a finite presence of gas. While the high field surfaces become extremely clean due to the action of the applied fields, the rest of the vacuum system surface has the adsorbed gas associated with the base vacuum pressure of the system. During an idle period, gas is transferred from the non high-field regions to the high-field ones. Extrinsic-associated vacuum and breakdown activity occur once fields are applied again. This is the same effect which is proposed to explain the dependency of breakdown rate on repetition rate and pause time that will be discussed subsection III.5.
@@
A few important insights can be made from these observations from conditioning. One is that there is not a unique type of breakdown trigger, but rather there are both those associated with extrinsic surface contaminants and those associated with intrinsic features coming from the fundamental properties of the device material surface or bulk. Long-term conditioning is determined by intrinsic material properties and takes many hundreds of millions of pulses and is preserved even when a device is exposed to air. Because the dislocation model predicts the observed field and temperature dependencies of breakdown rate, one may suppose that an evolving dislocation structure underlies this long-term conditioning. However differences in dislocation structure between conditioned and unconditioned areas has not yet been directly observed.
@@
Different materials, and materials in different metallurgical states condition in different ways and to different levels. For example, hard copper conditions faster than soft copper as shown in Fig. 20. This supports that long-term conditioning is driven by intrinsic properties of the system. We will return to dependence on material in subsection III.6.
@@
A crucial question for optimizing a conditioning strategy is if the conditioning speed is dependent on the level of applied field. As we have seen conditioning is correlated to the number of pulses as shown in Fig. 19 but the overall rate may be higher with applied field. Evidence that conditioning proceeds more quickly with higher applied field for dc systems is reported in [86], with key evidence reproduced here in Fig. 21. The anodes and cathodes in this test had the same diameter, which resulted in a field enhancement at the outer edge of about 7.5%. For the hard electrode, on the left in Fig. 21, breakdowns populate the outer edge because the hard electrode is in a more conditioned state, as can be seen in the faster conditioning shown in Fig. 20. On the other hand, breakdowns populate the soft electrode (on the right in Fig. 21) quite uniformly and are not concentrated on the outer edge despite the field enhancement. The explanation for this is that the higher field on the outer edges results in a higher conditioning speed compared to the center of the electrode, which offsets the strong dependency of breakdown rate on applied field.
@@
Taking this observation on the interplay between conditioning speed and breakdown rate through applied field further, and in order to better determine the mechanisms involved in conditioning a Monte Carlo simulation tool has been developed [123], also inspired by previous work [102] . The dual role of the applied field in both driving conditioning and the breakdown rate is a key element of understanding the breakdown process.